Please also include the correct error and how you calculated it.

A binary mixture with a total mass of 0.810 ± 0.001 g of BaCl2•2H2O (molar mass = 244.27 g/mol) and Na2SO4 (molar mass = 142.04 g/mol) produces 0.466 ± 0.001 g of BaSO4 (molar mass = 233.39 g/mol); BaCl2•2H2O is the limiting reactant. The total mass of 0.810 ± 0.001 g contains both of the salts in unknown amounts. Calculate the percent of BaCl2•2H2O and the percent of Na2SO4 in the original mixture with error.
Answer #1

Answer

Reaction between BaCl2.2H2O and Na2SO4 taking place as follows-

BaCl2.2H2O + Na2SO4 = BaSO4 + 2NaCl + 2H2O

Weight of BaSO4 produced = 0.466 \pm 0.001 g

Moles of BaSO4 produced = 0.466/142.04 = 0.00199 moles.

So, in the above reaction, 0.00199 moles of BaCl2.2H2O react with 0.00199 moles of Na2SO4 to produce 0.00199 moles of BaSO4.

So, weight of BaCl2.2H2O in the mixture = 0.00199 x 244.27 = 0.4860973 g

weight of Na2SO4 in the mixture = 0.00199 x 142.04 = 0.2826596 g

So, the percent of BaCl2.2H2O in the mixture = (0.4860973 x 100)/0.810 = 60.01 \pm 0.001 %

the percent of Na2SO4 in the mixture = (0.2826596 x 100)/0.810 = 34.89 \pm 0.001 %